OpenStudy (anonymous):

what are the solutions for x in the proportion 4x-3 over x + equal to 2x over 2

OpenStudy (anonymous):

\[\frac{4x - 3}{x} = \frac{2x}{2}\] Is that the problem?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Just cross multiply. \[2x(x) = 2(4x - 3)\] \[2x^{2} = 8x - 6\] \[2x^{2} - 8x + 6 = 0\] \[2(x^{2} - 4x + 3) = 0\] Can you solve the quadratic?

OpenStudy (anonymous):

I don't know how

OpenStudy (anonymous):

\[2(x^{2} - 4x + 3) = 0\] \[2(x - 3)(x - 1) = 0\] Can you solve for x now?